OpenStudy (anonymous):

Which factor can you cancel from the numerator and denominator to simplify (4x-16)/(12-3x) ? [Hint: (a−x) = − (x − a)]

OpenStudy (anonymous):

4 and 3

OpenStudy (mr.math):

\[\frac{4(x-4)}{-3(x-4)}=-\frac{4}{3}.\]

OpenStudy (mr.math):

So you cancel x-4.

OpenStudy (anonymous):

okay thank you :)

OpenStudy (mr.math):

You're welcome! :)